OpenStudy (anonymous):

If y=2x^2-4x-5 what is the value of y when x=-3?

OpenStudy (izzy18970):

i hope this helps

OpenStudy (anonymous):

you just replace the x in the given y :) like this y= 2 times ( -3)^2 -4 times (-3) -5 y = 2 times 9 +12 -5 y= 18+12-5 y= 25 And there you go, that basically all you need to do. hope this help

OpenStudy (anonymous):

Thank you. I also am stuck on this: Equivalent. 5r-3>2r+6? r>1/3. r>3 r>1. r>6

OpenStudy (anonymous):

Leong?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it just like this: 5r-3>2r+6? 5r-2r-3>2r-2r+6 3r-3>6 3r-3+3>6+3 3r>9 ( divide 3 for each side) r>3 So your answers should be r>3 :)